Solve each equation for x
x^2+4x+4=7
thanks

Answer by checkley75(3666)

X^2+4X+4=7
X^2+4X+4-7=0
X^2+4X-3=0
USING THE QUADRATIC EQUATION WE GET:
X=(-4+-SQRT[4^2-4*1*-3])/2*1
X=(-4+-SQRT[16+12])/2
X=(-4+-SQRT28])/2
X=(-4+-5.29)/2
X=(-4+5.29)/2
X=1.29/2
X=.645 ANSWER.
X=(-4-5.29)/2
X=-9.29/2
X=4.645 ANSWER.